OpenStudy (jholland):

does anyone know how to divide 10 divided by 4

OpenStudy (mertsj):

\[\frac{10}{4}=\frac{5}{2}=2 \frac{1}{2}=2.5\]

OpenStudy (anonymous):

it equals 2.5

OpenStudy (jholland):

Can u show me how to get that answer. I cant use calculators in school and I have no idea how to do this type of division thanks so very much